# Quantitative LLM Judges

### v1

@@ 4.2 Multinomial Judge
The multinomial (MN) judge is an absolute judge designed for predicting categorical scores, such as Likert scores. The Likert score can be viewed as an absolute score or as a ranking among options [6]. The MN judge is designed for the former interpretation. Our relative judges in Sections 4.3 and 4.4 are designed for the latter.
The MN judge predicts the most likely score from a set of scores $\mathcal{S}$. The probability of score $s\in\mathcal{S}$ is computed as
@@
where $\phi(e)\in\mathbb{R}^{d}$ is the embedding of base judge’s textual evaluation, $p_{s}$ is the probability that the base judge predicts score $s$, $p=(p_{s})_{s\in\mathcal{S}}\in\Delta^{\mathcal{S}}$ is a vector of all the probabilities, $\theta_{s}\in\mathbb{R}^{d+1}$ is a learned parameter for score $s$, and $c_{s}$ is a population bias towards score $s$. We denote all learned parameters by $\Theta=(\theta_{s})_{s\in\mathcal{S}}$ and estimate $p_{s}$ using the next token probability in $e$ [13, Section 2.3.3].
Since (2) is equivalent to the probability of outcome $s$ in multinomial logistic regression [29, Chapter 8], the population bias $c_{s}$ plays the same role. We introduce $p$ so that we can always learn a judge that performs at least as well as the base judge. Specifically, when $\theta_{s}=\mathbf{0}_{d}\oplus 1$ and $c_{s}=0$ for all $s\in\mathcal{S}$, the predicted probability becomes the base judge’s probability because
@@
We learn the judge by maximizing the probability of correct score predictions. This is equivalent to minimizing a regularized cross-entropy loss $\mathcal{L}(\theta)=-\sum_{t=1}^{n}\log\pi(s_{t}\mid e_{t},p_{t};\Theta)+\gamma\|\theta\|_{2}^{2}$, where $e_{t}$ is the textual evaluation, $p_{t}$ is a distribution over base judge’s scores, and $s_{t}$ is the human score in data point $t\in[n]$. The regularization strength $\gamma>0$ is set automatically by cross-validation.
We consider both LS and MN judges because they provide different perspectives on predicting an absolute score: regression versus classification. The LS judge treats the scores as real numbers and minimizes the squared error. The MN judge treats the scores as discrete choices and maximizes the accuracy of predicting them.
Absolute judgeRelative judgeAbsolute judgeAbsolute judgePrompt $x$Response $y$Prompt $x$Response $y_{1}$esponse $y_{2}$rompt $x$Response $y_{1}$esponse $y_{2}$ / MN judgeBTL judgeBTL2 judgeText $e$Score $b$Text $e$Score $b$Text $e_{1}$core $b_{1}$ext $e_{2}$core $b_{2}$bsolute scoreRelative scoreRelative score
Figure 1: Visualization of the architectures of the least-squares (LS), multinomial (MN), Bradley-Terry-Luce (BTL), and two-headed BTL (BTL2) judges described in Section 4.
@@ 4.3 Bradley-Terry-Luce Judge
The Bradley-Terry-Luce (BTL) judge is a relative judge that estimates the preference of one response over another from its evaluation by a relative base judge. The judge is motivated by arguably the most popular discrete choice model in human preference modeling [5]. The probability that the first response is preferred is computed as
@@
where $\mu$ is a sigmoid function, $\phi(e)\in\mathbb{R}^{d}$ is the embedding of base judge’s textual evaluation, $p$ is the probability that the base judge prefers the first response, $\theta\in\mathbb{R}^{d+1}$ is a learned model parameter, and $c$ is a population bias. We estimate $p$ using the next token probability in $e$ [13, Section 2.3.3]. The first response is preferred when $\pi(e,p;\theta)>0.5$; otherwise the second one is preferred.
@@
We learn the judge by maximizing the probability of ranking correctly. We pose this as minimizing a regularized logistic loss $\mathcal{L}(\theta)=-\sum_{t=1}^{n}[s_{t}\log\pi(e_{t},p_{t};\theta)+(1-s_{t})\log(1-\pi(e_{t},p_{t};\theta))]+\gamma\|\theta\|_{2}^{2}$, where $e_{t}$ is the textual evaluation, $p_{t}$ is the probability that the base judge prefers the first response, and $s_{t}\in\left\{0,1\right\}$ is the human score in data point $t\in[n]$. When $s_{t}=1$, the human prefers the first response; and when $s_{t}=0$, the human prefers the second one. The regularization strength $\gamma>0$ is set automatically by cross-validation.
@@ 4.4 Two-Headed BTL Judge
The two-headed BTL (BTL2) judge is a BTL judge that estimates the preferred response from two separate absolute evaluations. This builds on the findings that pointwise evaluators tend to be more robust [17], while pairwise evaluators are more susceptible to superficial cues due to inherent biases in LLMs [42, 8]. Our empirical results in Table 2 strongly support this approach.
We instantiate BTL2 within the framework of Section 4.3 as follows. Let $\phi(e_{1}),\phi(e_{2})\in\mathbb{R}^{d}$ be the embeddings of base judge’s textual evaluations $e_{1}$ and $e_{2}$, respectively. Let $b_{1},b_{2}>0$ be the base judge’s scores for $e_{1}$ and $e_{2}$, respectively. The probability that the first response is preferred is (3), where $\phi(e)=\phi(e_{1})-\phi(e_{2})$ and $p=b_{1}/(b_{1}+b_{2})$. We motivate this representation by noting that the difference of the embeddings $\phi(e)=\phi(e_{1})-\phi(e_{2})$ captures the difference of word affinities in the two responses. The probability $p$ biases the response based on base judge’s scores. Specifically, when $b_{1}>b_{2}$, $\log\frac{p}{1-p}>0$ and the first response is preferred; otherwise the second response is preferred. The judge is learned as in Section 4.3. We set $p_{t}=b_{t,1}/(b_{t,1}+b_{t,2})$, where $b_{t,1},b_{t,2}>0$ are the base judge’s scores for both responses in data point $t\in[n]$.
@@ 5 Experiments
To validate the performance of our quantitative judges, we comprehensively evaluated them on four tasks: two for absolute rating and two for relative preference prediction. We compare them to base judges as well as fine-tuned judges. We use $6$ metrics in rating tasks and $7$ in preference tasks. The metrics evaluate prediction errors, classification accuracy, and correlation.
@@ 5.1 Datasets and Metrics
Our datasets span both absolute rating and relative preference tasks.
[…]
HelpSteer2 [44, 45] is a dataset of absolute ratings for instruction-response pairs with correctness, coherence, complexity, verbosity, and overall helpfulness scores. The helpfulness is a score on a $5$-point Likert scale and we use it in our experiments. We train on their training set (20.3k data points) and test on their validation set (1.04k data points).
Offset Bias [32] is a synthetic pairwise preference dataset composed of instruction-response triplets $(x,y_{1},y_{2})$, where $x$ is a prompt, $y_{1}$ is a good response, and $y_{2}$ is a high-quality flawed response. This dataset is designed to confuse judges by injecting critical errors into otherwise fluent outputs, and is constructed using GPT-4 and Claude-3 with prompting strategies targeting off-topic and erroneous behavior. We create our own training (6.8k data points) and test (1.7k data points) sets from their publicly available training set.
Nectar [55] is a large-scale preference dataset where GPT-4 ranks responses from seven different models. We convert each data point into $\binom{7}{2}$ pairwise comparisons required by our BTL and BTL2 judges. Our own training (83.9k data points) and test (21k data points) sets are created from their publicly available training set.
@@
Summarize from Feedback and HelpSteer2 provide absolute ratings for language quality, covering summarization and instruction-following domains. They are widely adopted in the LLM evaluation literature. Offset Bias and Nectar are synthetic datasets that focus on relative comparison, testing robustness to adversarial scenarios and scalability. Our selection of datasets ensures that we test on both calibrated human judgments and challenging synthetic ground-truth.
@@
We consider three types of metrics. For rating prediction tasks, we report the mean squared error (MSE), mean absolute error (MAE), and Likert scale accuracy. For preference prediction tasks, we report accuracy (probability that the judge agrees with the ground-truth preference), precision $\Big{(}\frac{TP}{TP+FP}\Big{)}$, recall $\Big{(}\frac{TP}{TP+FN}\Big{)}$, and the F1-Score. TP, FP, and FN denote the number of true positives, false positives, and false negatives, respectively. We also report three correlation metrics: Pearson correlation coefficient ($r$), Spearman’s rank correlation coefficient ($\rho$), and Kendall’s $\tau$. The correlation metrics show the utility of the predictive scores. In particular, even if the scores cannot be predicted well, they may be predicted well enough for ranking them, which often suffices in practice.
@@
Naive 2.871 1.391 0.187 1.557 0.985 0.300
Table 1: Evaluation on rating prediction tasks. We report three prediction metrics (MSE, MAE, and accuracy) and three correlation metrics (Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$). The best result for each dataset and base judge is reported in bold. The underlined numbers are statistically significant gains over the absolute base judge at $p<0.05$. The naive baselines are the mean and majority Likert ratings.
@@
Table 2: Evaluation on preference prediction tasks. We report four prediction metrics (accuracy, precision, recall, and F-score) and three correlation metrics (Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$). The best result for each dataset and base judge is reported in bold. The underlined numbers are statistically significant gains over the best base judge at $p<0.05$.
5.2 Implementation Details
We have two types of baselines: the absolute and relative base judges described in Section 3, and their fine-tuned versions (SFT). We fine-tune the judges to predict human scores from their prompts instantiated with evaluated responses. We experiment with two base judges: Prometheus-7B-V2 [21] (Prometheus) and Llama-3.1-8B-Instruct [12] (Llama). Prometheus is a specialized model fine-tuned for evaluation whereas Llama is a general instruction-following model. We experiment with Llama to test if our framework generalizes beyond specialized evaluators.
Our evaluation protocol follows the grading methodology of Prometheus [21], which defines the two types of judges in Section 3. To ensure consistency across all judges, we use standardized grading prompts (Appendix A). All models are queried using vLLM and the same decoding configuration: temperature = 0.1, top_p = 0.9, and top_k = -1 (unrestricted sampling). Embeddings in our judges $\phi(e)$ are obtained from the base judge using the output of the final layer. Our judges are trained by SGD [35] and we set the regularization strength $\gamma$ by $5$-fold cross-validation.
All reported results are on test sets. Unless explicitly stated, we train on $10\%$ of our training sets, to mimic a real-world setting where the number of human ratings is from mid hundreds (HelpSteer2) to multiple thousands (Nectar). We show how the performance of our judges varies with training set size in Section 5.5 and Appendix B. All results are averaged over $10$ randomly chosen training sets. We use the multiple runs to determine statistical significance.
5.3 Results
Our results on absolute rating prediction tasks are reported in Table 1. We start with the LS judge on Summarize from Feedback dataset with the Prometheus base judge. The MSE of the LS judge ($2.626$) is less than a half of that of the base judge ($6.346$), and the lowest of all methods. This is expected because we optimize the MSE (Section 4.1). The LS judge has also the lowest MAE. The accuracy of the MN judge ($0.229$) is $36\%$ higher than that of the base judge ($0.168$). The MN judge is outperformed by SFT, which also maximizes the probability of predicting the correct score but using supervised fine-tuning (Section 5.2). The MN judge is trained in a fraction of the time of SFT (Section 5.4) though. Finally, although we optimize for rating prediction, the correlation metrics are comparable to the base judge, except for a drop in Kendall’s $\tau$. We report confusion matrices of our judges in Appendix C. We observe similar trends for LS and MN judges with the Llama base judge, except that the MN judge has the highest accuracy and all correlation metrics drop.
On HelpSteer2 dataset, with both Prometheus and Llama base judges, there are two changes in the trends. First, the LS judge significantly improves over the correlation metrics of the base judges. Second, SFT attains the lowest MAE with the Prometheus base judge. In summary, Table 1 shows that quantitative judges, despite their simplicity, outperform pre-trained judges on optimized metrics, and can outperform more complex and computationally costly alternatives.
Our results in Table 1 show that optimization of absolute scores can degrade ranking metrics. We optimize the ranking metrics in Table 2. On Offset Bias dataset with the Prometheus base judge, the BTL judge outperforms the relative base judge in all metrics. While the absolute judge outperforms the BTL judge, it is also bested by the BTL2 judge in all metrics. Notably, the Pearson correlation coefficient $r$ and Spearman’s rank correlation coefficient $\rho$ double when compared to the base judge. We observe the same major improvements for BTL and BTL2 judges with the Llama base judge. The results on Nectar dataset are less conclusive. While the BTL2 judge performs comparably to SFT with the Prometheus base judge, SFT performs the best with the Llama base judge. We note that the BTL and BTL2 judges can be trained in a fraction of the time for SFT (Section 5.4). In summary, quantitative judges can outperform pre-trained judges and more computationally costly alternatives.
Judge Summarize from Feedback HelpSteer2 Offset Bias Nectar
LS $0.320+0.455$ $0.880+0.450$
MN $0.320+0.468$ $0.880+0.461$
BTL - - $1.569+0.457$ $21.483+1.249$
BTL2 - - $2.247+0.538$ $34.731+1.221$
SFT $14.160$ $27.460$ $19.300$ $276.700$
Table 3: Training times in GPU minutes per dataset with Prometheus base judge. We break the times of our judges into computing embeddings plus training the judge.
5.4 Computation Times
Our quantitative judges have both inference and training components (Section 4). The computational overhead at inference time is minimal when implemented correctly, because the base judge’s embedding $\phi(e)$ is available for free after the evaluation $e$ is generated; and the score $b$ or its probability $p$ can be obtained in $O(1)$ time as $e$ is generated. Because of this, we only focus on training times in this discussion.
We report the total GPU training times of our quantitative judges and SFT in Table 3. The times are measured on an NVIDIA-A100-SXM4-80GB GPU. We observe two major trends. First, the times of our judges are often dominated by that of computing the base judge’s embeddings $\phi(e)$. Second, our times may be an order of magnitude smaller than for SFT. Notably, our training time on Offset Bias dataset is $19.3/(2.247+0.538)=6.93$ times lower and yet we outperform SFT in all metrics in Table 2. This is a direct benefit of the statistical and computational efficiency of simpler models at smaller sample sizes.
5.5 Ablation Studies

### v2

@@ 4.2 Multinomial Judge
The multinomial (MN) judge is an absolute judge designed for predicting categorical scores, such as on the Likert scale. The Likert score can be viewed as an absolute score or as a ranking among options (Carifio and Perla, 2008). The MN judge is designed for the former. Our relative judges in Sections 4.3 and 4.4 are designed for the latter.
The MN judge predicts the most likely score from a set $\mathcal{S}$. The probability of score $s\in\mathcal{S}$ is
@@
where $\phi(e)\in\mathbb{R}^{d}$ is the embedding of base judge’s rationale, $p_{s}$ is the probability that the base judge predicts score $s$, $p=(p_{s})_{s\in\mathcal{S}}\in\Delta^{\mathcal{S}}$ is a probability vector, $\theta_{s}\in\mathbb{R}^{d+1}$ is a learned model parameter for score $s$, and $c_{s}$ is a population bias towards $s$. We denote all learned parameters by $\Theta=(\theta_{s})_{s\in\mathcal{S}}$ and estimate $p_{s}$ using the next token probability in $e$ (Gu et al., 2025, Section 2.3.3).
Since (2) is equivalent to the probability of outcome $s$ in multinomial logistic regression (Murphy, 2012, Chapter 8), the population bias $c_{s}$ plays the same role. We introduce $p$ so that we can always learn a judge that performs at least as well as the base judge. Specifically, when $\theta_{s}=\mathbf{0}_{d}\oplus 1$ and $c_{s}=0$ for all $s\in\mathcal{S}$, the predicted probability becomes the base judge’s probability because
@@
We learn the judge by maximizing the probability of correct score predictions. This is equivalent to minimizing a regularized cross-entropy loss $\mathcal{L}(\Theta)=-\sum_{t=1}^{n}\log\pi(s_{t}\mid e_{t},p_{t};\Theta)+\gamma\|\Theta\|_{2}^{2}$, where $e_{t}$ is the base judge’s rationale, $p_{t}$ is a distribution over its scores, and $s_{t}$ is the human score in data point $t\in[n]$. The regularization strength $\gamma>0$ is set automatically by cross-validation.
We consider both LS and MN judges because they provide two different perspectives on predicting an absolute score: regression versus classification. The LS judge treats the scores as real numbers and minimizes the squared error. The MN judge treats the scores as discrete choices and maximizes the accuracy of predicting them.
@@ 4.3 Bradley-Terry-Luce Judge
The Bradley-Terry-Luce (BTL) judge is a relative judge that estimates the preference of one response over another from its evaluation by a relative base judge. The judge is motivated by the most popular discrete choice model in human preference modeling (Bradley and Terry, 1952). The probability that the first response is preferred is computed as
@@
where $\mu$ denotes a sigmoid function, $\phi(e)\in\mathbb{R}^{d}$ is the embedding of base judge’s rationale, $p$ is the probability that the base judge prefers the first response, $\theta\in\mathbb{R}^{d+1}$ is a learned model parameter, and $c$ is a population bias. We estimate $p$ using the next token probability in $e$ (Gu et al., 2025, Section 2.3.3). The first response is preferred when $\pi(e,p;\theta)>0.5$; otherwise the second one is preferred.
@@
We learn the judge by maximizing the probability of ranking correctly. We pose this as minimizing a regularized logistic loss $\mathcal{L}(\theta)=-\sum_{t=1}^{n}[s_{t}\log\pi(e_{t},p_{t};\theta)+(1-s_{t})\log(1-\pi(e_{t},p_{t};\theta))]+\gamma\|\theta\|_{2}^{2}$, where $e_{t}$ is the base judge’s rationale, $p_{t}$ is the probability that the judge prefers the first response, and $s_{t}\in\left\{0,1\right\}$ is the human score in data point $t\in[n]$. When $s_{t}=1$, the human prefers the first response; and when $s_{t}=0$, the human prefers the second one. The regularization strength $\gamma>0$ is set automatically by cross-validation.
@@ 4.4 Two-Headed BTL Judge
The two-headed BTL (BTL2) judge is a BTL judge that estimates the preferred response from two separate absolute evaluations. This builds on the findings that pointwise evaluators tend to be more robust (Jeong et al., 2024), while pairwise evaluators are more susceptible to superficial cues due to inherent biases in LLMs (Wang et al., 2024a; Chiang and Lee, 2023). Our empirical results in Table 2 strongly support this approach.
We instantiate BTL2 within the framework of Section 4.3 as follows. Let $\phi(e_{1}),\phi(e_{2})\in\mathbb{R}^{d}$ be the embeddings of base judge’s rationales $e_{1}$ and $e_{2}$, respectively; and $b_{1},b_{2}>0$ be the associated scores. We define the probability that the first response is preferred as (3), where $\phi(e)=\phi(e_{1})-\phi(e_{2})$ and $p=b_{1}/(b_{1}+b_{2})$. This representation is motivated by the fact that the difference of the embeddings $\phi(e)=\phi(e_{1})-\phi(e_{2})$ reflects the difference of word affinities in the two responses. The probability $p$ biases the response towards the base judge. Specifically, when $b_{1}>b_{2}$, $\log\frac{p}{1-p}>0$ and the first response is preferred; otherwise the second response is preferred. The judge is learned exactly as in Section 4.3. We set $p_{t}=b_{t,1}/(b_{t,1}+b_{t,2})$, where $b_{t,1},b_{t,2}>0$ are the base judge’s scores for both responses in data point $t\in[n]$. We extend BTL2 beyond a binary comparison in Section D.1.
@@ 5 Experiments
To validate the performance of our quantitative judges, we comprehensively evaluate them on four tasks: two for absolute rating and two for relative preference prediction. We compare them to their base judges, fine-tuning them, an order of magnitude larger judge, and specialized judges.
@@ 5.1 Datasets and Metrics
We experiment with $4$ datasets. The first two provide absolute ratings for language quality and are widely adopted in the LLM evaluation literature. The last two are synthetic datasets that focus on relative comparison, testing robustness to adversarial scenarios and scalability. Our dataset selection ensures that we test on both calibrated human judgments and challenging synthetic ground truth. We provide more details next.
Summarize from Feedback (Stiennon et al., 2020) is a human-annotated dataset with summary responses rated on a $7$-point Likert scale. We use its axis subset, which contains absolute scores for overall helpfulness, accuracy, coverage, and coherence. We use the overall score in our experiments, train on their validation set (8.598.59k data points), and test on their test set (6.316.31k data points).
HelpSteer2 (Wang et al., 2024c; d) is a dataset of absolute ratings for instruction-response pairs with correctness, coherence, complexity, verbosity, and overall helpfulness scores. The helpfulness is a score on a $5$-point Likert scale and we use it in our experiments. We train on their training set (20.320.3k data points) and test on their validation set (1.041.04k data points).
Offset Bias (Park et al., 2024) is a synthetic pairwise preference dataset composed of instruction-response triplets $(x,y_{1},y_{2})$, where $x$ is a prompt, $y_{1}$ is a good response, and $y_{2}$ is a high-quality flawed response. This dataset is designed to confuse judges by injecting critical errors into otherwise fluent outputs, targeting off-topic and erroneous behavior. We create our own training (6.86.8k data points) and test (1.71.7k data points) sets from their publicly available training set.
Nectar (Zhu et al., 2023) is a large-scale preference dataset where GPT-4 ranks responses from seven different models. We convert each data point into $\binom{7}{2}$ pairwise comparisons required by our BTL and BTL2 judges. We create our own training (83.983.9k data points) and test (2121k data points) sets from their publicly available training set.
@@
We consider three types of metrics. For rating prediction tasks, we report the mean squared error (MSE), mean absolute error (MAE), and accuracy. For preference prediction tasks, we report accuracy (probability that the judge agrees with the ground-truth preference), precision $\Big(\frac{TP}{TP+FP}\Big)$, recall $\Big(\frac{TP}{TP+FN}\Big)$, and the F1 score, where TP, FP, and FN are the number of true positives, false positives, and false negatives, respectively. In addition, we report three correlation metrics: Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$. The correlation metrics show the utility of the predicted scores, if they can be used for ranking responses.
5.2 Implementation Details
We experiment with two base judges: Llama-3.1-8B-Instruct (Grattafiori et al., 2024) (Llama) and Prometheus-7B-V2 (Kim et al., 2024b) (Prometheus). Prometheus is the most popular fine-tuned model for evaluation. We choose Llama because it is a popular generic model that can follow the same evaluation instructions as Prometheus. We employ both models as both absolute and relative judges (Section 3). The judges are implemented using the prompts in Appendix A, which we borrow from Kim et al. (2024b). Both prompts ask the model to reason and then output a score. Embeddings in our judges $\phi(e)$ are obtained from the base judge’s final layer. The regularization strength $\gamma$ in the quantitative judges is set by $5$-fold cross-validation and we use SGD (Robbins and Monro, 1951) to learn them. Unless explicitly stated, we train on $10\%$ of training sets to mimic a real-world setting where the number of human ratings is relatively small: from hundreds (Offset Bias) to thousands (Nectar). All results are averaged over $10$ random training sets to avoid bias to a given split.
Beyond the base judges, we consider three baselines. The first baseline is a supervised fine-tuned base judge to human scores from prompts instantiated with evaluated responses. We call it SFT. The second baseline is Llama-3.1-70B-Instruct as a base judge and we call it 70B. The last baseline is another specialized judge JudgeLM (Zhu et al., 2025). These baselines cover more computationally costly alternatives as well as another specialized judge. To have a fair comparison, all models are queried using vLLM with the same decoding configuration: $\textsf{temperature}=0.1$, $\textsf{top\_p}=0.9$, and $\textsf{top\_k}=-1$ (unrestricted sampling). All reported metrics are on test sets.
5.3 Results
@@
70B Absolute base 5.143 1.861 0.161 0.448 0.476 0.380 1.804 0.997 0.415 0.274 0.391 0.244
JudgeLM Absolute base 3.820 1.577 0.184 0.153 0.140 0.129 1.639 0.939 0.320 0.066 0.074 0.060
Table 1: Evaluation on rating prediction tasks. We report three prediction metrics (MSE, MAE, and accuracy) and three correlation metrics (Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$). The best result for each dataset and base judge is reported in bold. The underlined numbers are statistically significant gains over the corresponding absolute base judge at $p<0.05$.
Comparison to base judges and fine-tuning them: Our results on absolute rating prediction tasks are reported in Table 1. We start with Summarize from Feedback dataset and Prometheus base judge. The MSE of the LS judge ($2.626$) is less than a half of that of the base judge ($6.346$), and the lowest of all methods. This is expected because we optimize the MSE (Section 4.1). The LS judge has also the lowest MAE. The accuracy of the MN judge ($0.229$) is $36\%$ higher than that of the base judge ($0.168$). The MN judge is outperformed by SFT, which also maximizes the probability of predicting the correct score but using supervised fine-tuning (Section 5.2). Note that the training time of the MN judge is a fraction of that of SFT (Appendix C). Finally, although we optimize for rating prediction, the correlation metrics are similar to the base judge, except for a drop in Kendall’s $\tau$. We observe similar trends for LS and MN judges with Llama base judge, except that the MN judge has the highest accuracy and all correlation metrics drop. On HelpSteer2 dataset, with both Prometheus and Llama base judges, there are two changes. First, the LS judge significantly improves over the correlation metrics of the base judges. Second, SFT attains the lowest MAE with Prometheus base judge.
Our results on relative preference prediction tasks are reported in Table 2. On Offset Bias dataset with Prometheus base judge, the BTL judge outperforms the relative base judge in all metrics. While the absolute judge outperforms the BTL judge, the BTL2 judge bests it in all metrics. Notably, the Pearson’s $r$ and Spearman’s $\rho$ double when compared to the base judge. We observe the same major improvements for BTL and BTL2 judges with Llama base judge. The results on Nectar dataset are less conclusive. The BTL2 judge performs comparably to SFT when the base judge is Prometheus, but is outperformed when the base judge is Llama. Note that the training times of BTL and BTL2 judges are a fraction of that of SFT (Appendix C).
Our results in Tables 1 and 2 show that quantitative judges, despite their simplicity, can outperform pre-trained base judges on optimized metrics, as well as more computationally costly alternatives. The correlation metrics in Table 1 are much lower than in Table 2 because the quantitative judges in Table 1 are optimized to predict scores, not to rank them. The correlation metrics in Table 2 are high enough to be useful: a Kendall’s $\tau$ of $0.5$ corresponds to $75\%$ ranking accuracy when there are no ties. We attain a higher value than that in two experiments out of four.
Comparison to $10$-times larger judges: The absolute and relative variants of the 70B judge often outperform our original base judges because of a larger base model. Nevertheless, the quantitative judges remain mostly superior. In rating prediction tasks (Table 1), we improve over the 70B judge in MSE (LS judge), MAE (LS judge), and accuracy (MN judge). On Offset Bias dataset (Table 2), BTL2 significantly improves over the absolute 70B judge in $4$ metrics out of $7$ and is worse in $2$; and over the relative 70B judge in $6$ metrics out of $7$. On Nectar dataset (Table 2), BTL2 improves over the absolute 70B judge in all metrics. The relative 70B judge outperforms even SFT and is the best method on Nectar dataset.
Comparison to JudgeLM: Unlike the 70B judge, JudgeLM is not clearly better than our original base judges. In rating prediction tasks (Table 1), we improve over JudgeLM in MSE (LS judge), MAE (LS judge on Summarize from Feedback), and accuracy (MN judge). The correlation metrics of our judges are superior to those of JudgeLM. On Offset Bias dataset (Table 2), BTL2 significantly improves over both absolute and relative JudgeLM in all metrics. On Nectar dataset (Table 2), we observe the same trend with the absolute JudgeLM. The relative JudgeLM slightly outperforms BTL and both are ultimately outperformed by SFT.
@@
70B Absolute base 0.802 0.863 0.629 0.731 0.611 0.611 0.611 0.640 0.640 0.642 0.642 0.284 0.284 0.284
Relative base 0.680 0.811 0.468 0.593 0.397 0.397 0.397 0.784 0.793 0.753 0.772 0.569 0.569 0.569
JudgeLM Absolute base 0.585 0.484 0.167 0.248 0.060 0.060 0.060 0.568 0.414 0.277 0.332 0.032 0.032 0.032
Relative base 0.584 0.464 0.478 0.471 0.086 0.086 0.086 0.731 0.734 0.777 0.755 0.457 0.457 0.457
@@
Table 2: Evaluation on preference prediction tasks. We report four prediction metrics (accuracy, precision, recall, and F1 score) and three correlation metrics (Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$). The best result for each dataset and base judge is reported in bold. The underlined numbers are statistically significant gains over the best corresponding base judge at $p<0.05$.
5.4 Additional Studies
We conduct more experiments in Appendices B and C, and summarize their results here.
Human feedback: The amount of human feedback to learn from has a major impact on quantitative judges. By design, we can improve over base judges with very little data. As an example, in Tables 1 and 2, we learn from $10\%$ of training data, which is $859$ examples in Summarize from Feedback, $2030$ examples in HelpSteer2, $680$ examples in Offset Bias, and $8390$ examples in Nectar. We ablate how some metrics change with training data size in Figures 2, 3 and 4 (last two in Section B.1). The $1\%$ of training data in these figures corresponds to $86$ examples in Summarize from Feedback, $203$ examples in HelpSteer2, $68$ examples in Offset Bias, and $839$ examples in Nectar. Even at this operating point, our best judges (LS in Figure 2, MN in Figure 3, and BTL2 in Figure 4) outperform the best base judges in Tables 1 and 2. Note that when LLM judges are deployed, a small number of human scores is typically available because they are used to evaluate the judge. Our work gives a recipe for learning a better judge from these scores and provides empirical support for it.
Regularization: Automatic choice of the regularization strength $\gamma$ by cross-validation ensures that our judges perform well on test sets in Tables 1 and 2. Without it, our judges could perform poorly. For instance, Figure 5 in Section B.2 shows that the LS judge would have a higher MSE than the base judge when the regularization strength is high. On the other hand, the optimal regularization strength of the MN judge is dataset-dependent.
Embedding quality: Section B.3 studies the impact of embeddings $\phi(e)$ on quantitative judges. We conduct two experiments. First, we show that quantitative judges can be implemented with other embeddings than those of the base judge. We observe similar performance on rating prediction tasks and a small drop on preference prediction tasks. Second, we show that as the embeddings get worse, the quantitative judges get worse.
Alignment to human preferences: Quantitative judges predict scores from rationales and numeric scores of a frozen LLM. One way of measuring how the scores align with human preferences are the correlation metrics in Tables 1 and 2. In Table 1, the alignment may moderately improve or worsen, because we do not optimize it. We observe a major improvement in Table 2 though. On Offset Bias dataset with Llama base judge, the BTL2 judge doubles Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$ of the base judges. We delve deeper into how the improvements arise in Section B.4, where we report confusion matrices between human and predicted scores. In Figure 10, the base judge fails to predict frequent scores $5$ and $6$. Both the LS and MN judges learn to predict them, and in this way improve upon the base judge.
Computation time: We report the training and inference times of quantitative judges, and compare them to those of base judges and fine-tuning them, in Appendix C. The quantitative judges can be trained in an order of magnitude less time than supervised fine-tuning. The inference time overhead of the quantitative judges, dominated by computing the embedding of the base judge’s rationale, is about $25\%$ of that of generating the rationale; and thus not huge.